OpenStudy (anonymous):

Find a pair of factors for each number by using the difference of two squares. a. 45 b. 77 c. 112

OpenStudy (anonymous):

a. (5)(9) b. (7)(11) c. (8)(14) Does this look correct @gezimbasha Sorry to bother you

OpenStudy (anonymous):

Yes, they are factors. It is correct. But what's the thing with the difference of two squares?

OpenStudy (anonymous):

I dont know. maybe like this or not. a) 45 = 7^2 - 2^2 = (7 + 2)(7 - 2) = 9 * 5 b) 77 = 9^2 - 2^2 = (9 + 2)(9 - 2) = 11 * 7 c) 112 = 11^2 - 3^2 = (11 + 3)(11 - 3) = 14 * 8

OpenStudy (anonymous):

Yeah this one looks more like it. You're pretty good at this aren't you?!